# Unexpected properties of Markov chains with an infinitely countable set of states

### v4

@@ 2.3. Chapman-Kolmogorov equations
$+\mathsf{I}\{Z(t-)=n\}\mathsf{I}\{\mathrm{M}_{n}(t)-\mathrm{M}_{n}(t-)=0\}\mathsf{I}\{\tilde{\Lambda}_{n}(t)-\tilde{\Lambda}_{n}(t-)=0\},$
where $\mathrm{M}_{n}(t)$ denotes the Poisson process with rate $\mu_{n}$, $\tilde{\Lambda}_{n}(t)=2\Lambda_{n}(t)$, $\Lambda_{n}(t)$ denotes the Poisson process with rate $\lambda_{n}/2$.
For $n=0,1$, we have
(9) $\mathsf{I}\{Z(t)=1\}=\mathsf{I}\{Z(t-)=2\}\mathsf{I}\{\mathrm{M}_{2}(t)-\mathrm{M}_{2}(t-)=1\}$
@@
(10) $\mathsf{I}\{Z(t)=0\}=\mathsf{I}\{Z(t-)=1\}\mathsf{I}\{\mathrm{M}_{1}(t)-\mathrm{M}_{1}(t-)=1\}$
@@
From (8), (9) and (10) we arrive at the system of stochastic differential equations:
(11) $\mathrm{d}\mathsf{I}\{Z(t)=n\}=$ $\mathsf{I}\{Z(t-)=n+1\}\mathrm{d}\mathrm{M}_{n+1}(t)$
@@
Let us now derive (5) from (11). The derivation of (6) and (7) is similar and therefore not considered here. For the Poisson processes $\mathrm{M}_{n}(t)$ and $\Lambda_{n}(t)$ we use semimartingale decompositions $\mathrm{M}_{n}(t)=\mu_{n}t+m_{\mathrm{M}_{n}}(t)$ and $\Lambda_{n}(t)=(\lambda_{n}/2)t+m_{\Lambda_{n}}(t)$, where $m_{\mathrm{M}_{n}}(t)$ and $m_{\Lambda_{n}}(t)$ are corresponding martingales (see e.g. [5, 7, 9]). Then, taking into account that $\mathsf{E}m_{\mathrm{M}_{n}}(t)=\mathsf{E}m_{\Lambda_{n}}(t)=0$, from (11) we obtain:
$0=$ $\lim_{T\to\infty}\frac{1}{T}\mathsf{E}\int_{0}^{T}\mathsf{I}\{Z(t-)=n+1\}\mathrm{d}\mathrm{M}_{n+1}(t)$
$+\lim_{T\to\infty}\frac{1}{T}\mathsf{E}\int_{0}^{T}2\mathsf{I}\{Z(t-)=n-2\}\mathrm{d}\Lambda_{n-2}(t)$
$-\lim_{T\to\infty}\frac{1}{T}\mathsf{E}\int_{0}^{T}\mathsf{I}\{Z(t-)=n\}(\mathrm{d}\mathrm{M}_{n}(t)+2\mathrm{d}\Lambda_{n}(t))$
$=$ $\lim_{T\to\infty}\frac{1}{T}\mathsf{E}\int_{0}^{T}\mathsf{I}\{Z(t-)=n+1\}\mu_{n+1}\mathrm{d}t$
$+\lim_{T\to\infty}\frac{1}{T}\mathsf{E}\int_{0}^{T}\mathsf{I}\{Z(t-)=n-2\}\lambda_{n-2}\mathrm{d}t$
$-\lim_{T\to\infty}\frac{1}{T}\mathsf{E}\int_{0}^{T}\mathsf{I}\{Z(t-)=n\}(\mu_{n}+\lambda_{n})\mathrm{d}t$
$=$ $\mu_{n+1}Q_{n+1}+\lambda_{n-2}Q_{n-2}-Q_{n}(\lambda_{n}+\mu_{n}).$
Relation (5) is proved.
Indeed, from (3), (4) we have
$\lambda_{0}P_{0}+\sum_{n=1}^{\infty}P_{n}(\lambda_{n}+\mu_{n})=\mu_{1}P_{1}+\sum_{n=1}^{\infty}(\lambda_{n-1}P_{n-1}+\mu_{n+1}P_{n+1}),$
and from (5)–(7) we have
$=\mu_{1}Q_{1}+\sum_{n=1}^{\infty}(\lambda_{n-1}Q_{n-1}+\mu_{n+1}Q_{n+1}).$
Hence, $\sum_{n=0}^{\infty}Q_{n}=1$ if and only if $\sum_{n=0}^{\infty}P_{n}=1$.
Example 2.2.
Let $\mathscr{C}$ be a stochastic process with the same parameters $\lambda_{n}/2$ and $\mu_{n}$ as these in the birth-and-death process $\mathscr{B}$, but in any state $n$ there can be either one birth with probability $1/2$ or three births with the same probability $1/2$. It will be shown that the process $\mathscr{C}$ is recurrent if and only if the birth-and-death process $\mathscr{A}$ is recurrent.
Here we have the following system of difference equations
(12) $Q_{n}\left(\lambda_{n}+\mu_{n}\right)$ $=$ $\mu_{n+1}Q_{n+1}+\frac{\lambda_{n-1}}{4}Q_{n-1}+\frac{3\lambda_{n-3}}{4}Q_{n-3},\ n\geq 3,$
(13) $Q_{2}(\lambda_{2}+\mu_{2})$ $=$ $\mu_{3}Q_{3}+\frac{\lambda_{1}}{4}Q_{1},$
(14) $Q_{1}(\lambda_{1}+\mu_{1})$ $=$ $\mu_{2}Q_{2}+\frac{\lambda_{0}}{4}Q_{0},$
(15) $\lambda_{0}Q_{0}$ $=$ $\mu_{1}Q_{1}$
where $Q_{0}$, $Q_{1}$,…denote the final probabilities. The derivation of the system of equations (12)–(15) is similar to that given in Example 2.1 for the system of equations (5)–(7).
Similarly to the consideration given in Example 2.1, we arrive at
$\lambda_{0}Q_{0}+\sum_{n=1}^{\infty}(\lambda_{n}+\mu_{n})Q_{n}=\lambda_{0}P_{0}+\sum_{n=1}^{\infty}(\lambda_{n}+\mu_{n})P_{n},$
that means that $\sum_{n=0}^{\infty}Q_{n}=1$ if and only if $\sum_{n=0}^{\infty}P_{n}=1$. Hence, we arrive at the conclusion that the behavior of the stochastic process $\mathscr{C}$ is similar to the behavior of the birth-and-death process $\mathscr{A}$. Thus, all three processes $\mathscr{A}$, $\mathscr{B}$ and $\mathscr{C}$ are either recurrent or transient.
Apparently, the models considered in Examples 2.1 and 2.2 can be further extended, and from the point of view of recurrence or transience we will have the class of equivalent processes satisfying the same property. The class of equivalent processes should strongly contain the models of Markov chains that form connected domain only, although hitherto the provided proofs did not use any restriction. (The elementary models considered in Examples 2.1 and 2.2 do not form connected domain.)
Example 2.3.
$\left(\begin{matrix}0&1&0&0&0&0&0&0&0&\ldots\\ q_{1,1}&0&p_{1,1}&p_{1,2}&0&0&0&0&0&\ldots\\ q_{2,1}&q_{2,2}&0&p_{2,1}&p_{2,2}&p_{2,3}&0&0&0&\ldots\\ q_{3,1}&q_{3,2}&q_{3,3}&0&p_{3,1}&p_{3,2}&p_{3,3}&p_{3,4}&0&\ldots\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\ddots\end{matrix}\right),$
where all $p_{i,j}$ and $q_{i,j}$ that appear in the matrix are assumed to be strictly positive and $\sum_{j=1}^{i}q_{i,j}+\sum_{j=1}^{i+1}p_{i,j}=1$, $i=1,2,\ldots$, and the Markov chain of the birth-and-death type (call it $\mathscr{E}$), the matrix of transition probabilities of which is
$\left(\begin{matrix}0&1&0&0&0&0&\ldots\\ c_{1}q_{1,1}&0&c_{1}(p_{1,1}+2p_{1,2})&0&0&0&\ldots\\ 0&c_{2}(2q_{2,1}+q_{2,2})&0&c_{2}\sum_{j=1}^{3}jp_{2,j}&0&0&\ldots\\ 0&0&c_{3}\sum_{j=1}^{3}(4-j)q_{3,j}&0&c_{3}\sum_{j=1}^{4}jp_{3,j}&0&\ldots\\ \vdots&\vdots&\vdots&\ddots&\ddots&\ddots&\ddots\end{matrix}\right),$
where the normalization constants $c_{i}$ are given such that
$c_{i}\left(\sum_{j=1}^{i}(i+1-j)q_{i,j}+\sum_{j=1}^{i+1}jp_{i,j}\right)=1,\quad i=1,2,\ldots.$
(16) $c_{i}\left(\sum_{j=1}^{i}(i+1-j)q_{i,j}+\sum_{j=1}^{i+1}jp_{i,j}\right)P_{i}$
$=c_{i-1}P_{i-1}\sum_{j=1}^{i}jp_{i-1,j}+c_{i+1}P_{i+1}\sum_{j=1}^{i+1}(i+2-j)q_{i+1,j},$
(17) $P_{0}=c_{1}q_{1,1}P_{1},$
where $P_{0},P_{1},\ldots$ denote the final probabilities (if exist). In turn, for the Markov chain $\mathscr{D}$ we have the system of equation
(18) $c_{i}\left(\sum_{j=1}^{i}(i+1-j)q_{i,j}+\sum_{j=1}^{i+1}jp_{i,j}\right)Q_{i}$
$=c_{i-1}\sum_{j=1}^{i}jp_{i-1,j}Q_{i-j}+c_{i+1}\sum_{j=1}^{i+1}(i+2-j)q_{i+1,j}Q_{i+j},$
(19) $Q_{0}=c_{1}q_{1,1}Q_{1},$
where $Q_{0},Q_{1},\ldots$ denote the final probabilities (if exist). This last system of equations (18)–(19) is derived by the same method as it given in Example 2.1.
From these systems of equations (16)–(17) and (18)–(19) we respectively obtain
$P_{0}+\sum_{i=1}^{\infty}c_{i}\left(\sum_{j=1}^{i}(i+1-j)q_{i,j}+\sum_{j=1}^{i+1}jp_{i,j}\right)P_{i}$
$=c_{1}q_{1,1}P_{1}+\sum_{i=1}^{\infty}\left(c_{i-1}P_{i-1}\sum_{j=1}^{i}jp_{i-1,j}+c_{i+1}P_{i+1}\sum_{j=1}^{i+1}(i+2-j)q_{i+1,j}\right)$
and
$Q_{0}+\sum_{i=1}^{\infty}c_{i}\left(\sum_{j=1}^{i}(i+1-j)q_{i,j}+\sum_{j=1}^{i+1}jp_{i,j}\right)Q_{i}$
$=c_{1}q_{1,1}Q_{1}+\sum_{i=1}^{\infty}\left(c_{i-1}\sum_{j=1}^{i}jp_{i-1,j}Q_{i-j}+c_{i+1}\sum_{j=1}^{i+1}(i+2-j)q_{i+1,j}Q_{i+j}\right)$
$=c_{1}q_{1,1}Q_{1}+\sum_{i=1}^{\infty}\left(c_{i-1}Q_{i-1}\sum_{j=1}^{i}jp_{i-1,j}+c_{i+1}Q_{i+1}\sum_{j=1}^{i+1}(i+2-j)q_{i+1,j}\right).$
Hence,
$P_{0}+\sum_{i=1}^{\infty}c_{i}\left(\sum_{j=1}^{i}(i+1-j)q_{i,j}+\sum_{j=1}^{i+1}jp_{i,j}\right)P_{i}$
$=Q_{0}+\sum_{i=1}^{\infty}c_{i}\left(\sum_{j=1}^{i}(i+1-j)q_{i,j}+\sum_{j=1}^{i+1}jp_{i,j}\right)Q_{i},$
and as before we arrive at the conclusion that $\sum_{n=0}^{\infty}Q_{n}=1$ if and only if $\sum_{n=0}^{\infty}P_{n}=1$.
@@
Let us return to our proof. The proof of the theorem does not distinguish from the proof that has already been provided for the models considered in Examples 2.12.3. Despite the model considered in Example 2.3 is particular, all the features of that proof for the general model remain unchanged. Hence, the only thing that is required to do is to consider the associated birth-and-death process and derive the required system of equations.
@@
(20) $\mathsf{P}\{Z(t+\triangle t)=i\}$ $=\mathsf{P}\{Z(t+\triangle t)=i~{}|~{}A_{i}^{-}(\triangle t)\}e_{i+1}^{-}\triangle t$
@@
(21) $\mathsf{P}\{Z(t+\triangle t)=0\}$ $=\mathsf{P}\{Z(t+\triangle t)=0~{}|~{}A_{1}^{-}(\triangle t)\}e_{1}^{-}\triangle t$
@@
Here in (20) and (21), the meanings of the events $A_{i}^{-}(\triangle t)$ and $A_{i}^{+}(\triangle t)$ are that during time $(t,t+\triangle t]$ a down-crossing to state $i$ occurs and, respectively, an up-crossing to state $i$ occurs. The meanings of $e_{i}^{-}$ and $e_{i}^{+}$, $i\geq 1$, are the birth and, respectively, death rates in state $i$; $e_{0}^{+}=\sum_{j=1}^{\infty}jp_{0,j}$. Then the system of equations for the final probabilities of the birth-and-death process follows by the standard way. We have:
@@
3. Example
Let us return to Example 2.3 of Section 2.3. Assuming that
$q_{i,1}=q_{i,2}=\ldots=q_{i,i}\quad\text{and}\quad p_{i,1}=p_{i,2}=\ldots=p_{i,i+1},\quad i=1,2,\ldots,$
and
$\sum_{j=1}^{i}q_{i,j}=\sum_{j=1}^{i+1}p_{i,j}=\frac{1}{2},$
we obtain the following matrix of transition probabilities:
$\left(\begin{matrix}0&1&0&0&0&0&0&0&0&\ldots\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\frac{1}{2}&0&\frac{1}{4}&\frac{1}{4}&0&0&0&0&0&\ldots\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\frac{1}{4}&\frac{1}{4}&0&\frac{1}{6}&\frac{1}{6}&\frac{1}{6}&0&0&0&\ldots\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\frac{1}{6}&\frac{1}{6}&\frac{1}{6}&0&\frac{1}{8}&\frac{1}{8}&\frac{1}{8}&\frac{1}{8}&0&\ldots\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\ddots\end{matrix}\right).$
Our task is to find whether or not the Markov chain is recurrent. For associated birth-and-death process we have the following matrix of transition probabilities:
$\left(\begin{matrix}0&1&0&0&0&0&\ldots\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\frac{2}{5}&0&\frac{3}{5}&0&0&0&\ldots\\ \vskip 3.0pt plus 1.0pt minus 1.0pt0&\frac{3}{7}&0&\frac{4}{7}&0&0&\ldots\\ \vskip 3.0pt plus 1.0pt minus 1.0pt0&0&\frac{4}{9}&0&\frac{5}{9}&0&\ldots\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\ddots\end{matrix}\right)=\left(\begin{matrix}0&1&0&0&0&0&\ldots\\ \vskip 3.0pt plus 1.0pt minus 1.0ptq_{1}&0&p_{1}&0&0&0&\ldots\\ \vskip 3.0pt plus 1.0pt minus 1.0pt0&q_{2}&0&p_{2}&0&0&\ldots\\ \vskip 3.0pt plus 1.0pt minus 1.0pt0&0&q_{3}&0&p_{3}&0&\ldots\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\ddots\end{matrix}\right),$
where the matrix on the right-hand side is added for the notational convenience.
It is readily seen that the entries of this matrix satisfy the following property: ${p_{n}}/{q_{n}}=1+(n+1)^{-1}$ for all $n\geq 1$. Recall that the birth-and-death process is recurrent if and only if $\sum_{n=1}^{\infty}\prod_{i=1}^{n}q_{i}/p_{i}=\infty$. According to Raabe’s test, if ${p_{n}}/{q_{n}}\leq 1+n^{-1}$ for all large $n$, then the aforemention series diverges. In our case it is satisfied. Hence, the birth-and-death process is recurrent. Consequently, the Markov chain is recurrent.

### v5

@@ 2.3. Stochastic Chapman-Kolmogorov equations
$+\mathsf{I}\{Z(t-)=n\}\mathsf{I}\{\mathrm{M}_{n}(t)-\mathrm{M}_{n}(t-)=0\}\mathsf{I}\{\tilde{\Lambda}_{n}(t)-\tilde{\Lambda}_{n}(t-)=0\}$
that is valid with probability 1, where $\mathrm{M}_{n}(t)$ denotes the Poisson process with rate $\mu_{n}$, $\tilde{\Lambda}_{n}(t)=2\Lambda_{n}(t)$, $\Lambda_{n}(t)$ denotes the Poisson process with rate $\lambda_{n}/2$.
For $n=0,1$, with probability 1 we have
(2.37) $\mathsf{I}\{Z(t)=1\}=\mathsf{I}\{Z(t-)=2\}\mathsf{I}\{\mathrm{M}_{2}(t)-\mathrm{M}_{2}(t-)=1\}$
@@
(2.38) $\mathsf{I}\{Z(t)=0\}=\mathsf{I}\{Z(t-)=1\}\mathsf{I}\{\mathrm{M}_{1}(t)-\mathrm{M}_{1}(t-)=1\}$
@@
From (2.36), (2.37) and (2.38) we arrive at the system of stochastic differential equations, which is an analogue of (2.28):
$\mathrm{d}\mathsf{I}\{Z(t)=n\}=$ $\mathsf{I}\{Z(t-)=n+1\}\mathrm{d}\mathrm{M}_{n+1}(t)$
@@
The following arguments for the derivation of (2.6)–(2.8) are similar to that provided earlier for (2.17), (2.18).
Derivation of (2.11)–(2.14) is technically similar to the derivations of the systems of equations provided above.
@@
2.4. The final part of the proof
Let us return to our proof. The proof of the theorem does not distinguish from the proof that has already been provided for the model considered in Example 2.3. Despite the model considered in Example 2.3 is particular, the derivation of the system of the equations remains the same in general case. Hence, the only thing that is required to do is to consider the associated birth-and-death process and derive the required system of equations.
@@ 2.4. The final part of the proof
(2.39) $\mathsf{P}\{Z(t+\triangle t)=i\}$ $=\mathsf{P}\{Z(t+\triangle t)=i~{}|~{}A_{i}^{-}(\triangle t)\}e_{i+1}^{-}\triangle t$
@@
(2.40) $\mathsf{P}\{Z(t+\triangle t)=0\}$ $=\mathsf{P}\{Z(t+\triangle t)=0~{}|~{}A_{1}^{-}(\triangle t)\}e_{1}^{-}\triangle t$
@@
Here in (2.39) and (2.40), the meanings of the events $A_{i}^{-}(\triangle t)$ and $A_{i}^{+}(\triangle t)$ are that during time $(t,t+\triangle t]$ a down-crossing to state $i$ occurs and, respectively, an up-crossing to state $i$ occurs. The meanings of $e_{i}^{-}$ and $e_{i}^{+}$, $i\geq 1$, are the birth and, respectively, death rates in state $i$; $e_{0}^{+}=\sum_{j=1}^{\infty}jp_{0,j}$. Then the system of equations for the final probabilities of the birth-and-death process follows by the standard way. We have:
@@
3. Examples
In this section, we provide elementary but non-trivial examples of recurrent and transient Markov chains with an infinite set of states.
1. Consider the following matrix of transition probabilities
$\left(\begin{matrix}0&1&0&0&0&0&0&\cdots\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\frac{7}{12}&0&\frac{3}{12}&\frac{2}{12}&0&0&0&\cdots\\ \vskip 3.0pt plus 1.0pt minus 1.0pt0&\frac{7}{12}&0&\frac{3}{12}&\frac{2}{12}&0&0&\cdots\\ \vskip 3.0pt plus 1.0pt minus 1.0pt0&0&\frac{7}{12}&0&\frac{3}{12}&\frac{2}{12}&0&\cdots\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\vdots&\vdots&\vdots&\ddots&\ddots&\ddots&\ddots&\vdots\end{matrix}\right)$
The associated transition probability matrix of the birth-and-death type is
$\left(\begin{matrix}0&1&0&0&0&0&\cdots\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\frac{7}{10}&0&\frac{3}{10}&0&0&0&\cdots\\ \vskip 3.0pt plus 1.0pt minus 1.0pt0&\frac{1}{2}&0&\frac{1}{2}&0&0&\cdots\\ \vskip 3.0pt plus 1.0pt minus 1.0pt0&0&\frac{1}{2}&0&\frac{1}{2}&0&\cdots\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\vdots&\vdots&\vdots&\ddots&\ddots&\ddots&\vdots\end{matrix}\right)=\left(\begin{matrix}0&1&0&0&0&0&\cdots\\ \vskip 3.0pt plus 1.0pt minus 1.0ptq_{1}&0&p_{1}&0&0&0&\cdots\\ \vskip 3.0pt plus 1.0pt minus 1.0pt0&q_{2}&0&p_{2}&0&0&\cdots\\ \vskip 3.0pt plus 1.0pt minus 1.0pt0&0&q_{3}&0&p_{3}&0&\cdots\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\vdots&\vdots&\vdots&\ddots&\ddots&\ddots&\vdots\end{matrix}\right).$
With $q_{n}=p_{n}=1/2$ for $n\geq 2$, the birth-and-death process is recurrent. Hence, the above Markov chain is recurrent.
2. In the following example that is similar to the example above, the orders of fractions $3/12$ and $2/12$ in the matrix are changed:
$\left(\begin{matrix}0&1&0&0&0&0&0&\cdots\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\frac{7}{12}&0&\frac{2}{12}&\frac{3}{12}&0&0&0&\cdots\\ \vskip 3.0pt plus 1.0pt minus 1.0pt0&\frac{7}{12}&0&\frac{2}{12}&\frac{3}{12}&0&0&\cdots\\ \vskip 3.0pt plus 1.0pt minus 1.0pt0&0&\frac{7}{12}&0&\frac{2}{12}&\frac{3}{12}&0&\cdots\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\vdots&\vdots&\vdots&\ddots&\ddots&\ddots&\ddots&\vdots\end{matrix}\right)$
In this case, the Markov chain is transient.
Indeed, the associated transition probability matrix for the birth-and-death process now is
$\left(\begin{matrix}0&1&0&0&0&0&\cdots\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\frac{7}{9}&0&\frac{2}{9}&0&0&0&\cdots\\ \vskip 3.0pt plus 1.0pt minus 1.0pt0&\frac{7}{15}&0&\frac{8}{15}&0&0&\cdots\\ \vskip 3.0pt plus 1.0pt minus 1.0pt0&0&\frac{7}{15}&0&\frac{8}{15}&0&\cdots\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\vdots&\vdots&\vdots&\ddots&\ddots&\ddots&\vdots\end{matrix}\right)=\left(\begin{matrix}0&1&0&0&0&0&\cdots\\ \vskip 3.0pt plus 1.0pt minus 1.0ptq_{1}&0&p_{1}&0&0&0&\cdots\\ \vskip 3.0pt plus 1.0pt minus 1.0pt0&q_{2}&0&p_{2}&0&0&\cdots\\ \vskip 3.0pt plus 1.0pt minus 1.0pt0&0&q_{3}&0&p_{3}&0&\cdots\\ \vskip 3.0pt plus 1.0pt minus 1.0pt\vdots&\vdots&\vdots&\ddots&\ddots&\ddots&\vdots\end{matrix}\right).$
Now we have $p_{n}>q_{n}$ for all $n\geq 2$, and hence the transience of the birth-and-death process and the related Markov chain follows.